10. Jane is 54 tall, 25 years old and currently weighs 150 pounds. Her activity level is low- active (30 minutes per day)
Answer #1

Ans. 1) BMI is kg/m2 ​​​​​, where kg is weight in kilograms and m2 is height in meter square. Here height is given in feet, as 1 foot = 0.305 m, so 5'4 feet = 1.63 m. Similarly, weight is given in pounds, as 1 pound = 0.454 kg, so 150 pounds = 68 kg. So according to the formula, BMI = 68/ (1.63)2 = 25.66. A BMI of less than 18.5 means that a person is underweight, if between 18.5 and 24.9 is ideal, if between 25 and 29.9 is overweight and BMI over 30 indicates obesity. So here Jane's BMI is 25.6 that mean she is overweight, to be in range of ideal condition she need to lose minimum 4.4 pound to maximum 41.8 pound to bring BMI to 24.9 - 18.5 range.

Ans 2) for female above 19 year old, EER = 447.6 - 7.95*age (years) + PAL*(11.4* weight (kg) + 619* height (m)),

where PAL (physical activity level) = 1 if sedentary, 1.16 if low active, 1.27 if active and 1.44 if very active.

So on solving the above equation, EER = 2367.3 cal/day, it means she need this much kcal to maintain he weight. But as she is overweight she need to loss 1 to 1.5 pounds per week or approximately 6 pounds per month.   If she eliminate 500 kcal per day from her diet (or approximately 3500 kcal/week), she should be on track to meet this degree of weight loss. There is approximately 3500 calories per one pound of fat (0.45 kg). After subtracting 500 calories, new target (calories/day) would be 1867.3 which should result in roughly one-pound of weight loss per week.

Ans. 3) if her weight remain same and her activity increases her EER will be 2494.8 kcal/day. ( from the above equation).

Ans. 4) in higher activity level EER will be 2822.8 kcal/day, after subtracting 500 calories, new target (calories/day) would be 2322.8, which should be one pound loss per week and she need to lose minimum 4.4 pound so it would take 4-5 weeks i.e., approximately little more than one month to lose weight, if she maintain her higher activity per day.